# Learn from Heterophily: Heterophilous Information-enhanced Graph Neural Network

### v1

@@ 4.2 Performance
Table 3 shows a comparison of HiGNN with other baselines. All methods are evaluated with node classification task, where both the mean accuracy and standard deviation are reported based on 10 public random splits. The average rank for each method is calculated across 9 datasets, with a lower rank indicating better performance. For each dataset, the best-performing methods are highlighted in bold and the runner-up methods are underlined. For methods that augment general GNN baselines with heterophilous information (denoted as +Hi), we halve the embedding space to control the total number of parameters and time complexity, ensuring a fair comparison. If there is an improvement over the original methods, the results are displayed in blue.
@@
First, our proposed HiGNN generally outperforms all the other methods, achieving the lowest average rank, indicating the effectiveness of incorporating heterophilious information into graph learning. HiGNN enhances the performance on both the homophilous and heterophilous datasets, corroborating the explanation provided in Sect. 3.2 that the new matrix constructed by heterophilous information improves graph learning at different homophily levels. Besides, the effectiveness of heterophilous information is also shown in the improvement on the general GNNs baselines where the performances are improved on most of the datasets. Second, compared with general GNN baselines such as GCN and GAT, MLP performs poorly on homophilous graphs but still demonstrates strong performance on heterophilous graphs. This validates the assertion that heterophily impedes graph convolution in graph learning. Then for the methods that aggregate embeddings of k-hop neighbors during the message passing, Mixhop, H2GCN, GGCN, CPGCN, and GPR-GNN show their improvement with respect to GCN, but they still have a large gap with HiGNN and GloGNN that exploit global similar nodes. This indicates that for heterophilous graphs, focusing on k-hop neighbors is not as sufficient as constructing new graph structures that capture neighbors with similar semantics. Note that although GloGNN delivers the best results on 3 datasets, it requires the calibrated adjustment of many parameters. In contrast, our method is more straightforward and flexible, requiring fewer parameter tunings.
Table 3: Node classification performance of HiGNN
@@
4.3 Improvement in Homophily Degree
To validate that the new graph structure $A^{\prime}$ with heterophilous information in HiGNN improves homophily, we measure the homophily degree for both the original graph structure $A$ and the new graph structure $A^{\prime}$. Specifically, we compute homophily degrees across 9 datasets on 10 random splits and then report the mean and standard deviation. As depicted in Fig. 3, the newly constructed matrix improves the homophily degree on all datasets, particularly for heterophilous datasets. This suggests that heterophilous information is helpful in grouping semantically similar nodes. It is worth noting that the estimation of heterophilous information is based on the predicted labels instead of the true labels, demonstrating the estimation is error-tolerant as suggested in Sect. 3.3
@@ 4.3 Improvement in Homophily Degree
4.4 Ablation Study
To demonstrate the performance improvement of HiGNN achieved through the incorporation of heterophilous information and the efficacy of the fusion strategy employed in HiGNN, we undertake an ablation study. As shown in Table 4, for HiGNN, we examine the influence of channels by removing the new channel (w/o $A^{\prime}$), removing the original channel (w/o $A$), or fusing two adjacency matrices $A$ and $A^{\prime}$ in one channel ($A$+$A^{\prime}$). The results show that the performance of HiGNN drops when either the new channel or the original channel is removed. This suggests that optimal performance is attained by considering different aspects of these two channels. Furthermore, the early fusion of two adjacency matrices ($A$+$A^{\prime}$) underperforms compared to the late fusion in HiGNN, indicating the late fusion strategy allows the model to fully capture the unique contributions of each channel as suggested in Sect. 3.3.
Table 4: Ablation study on HiGNN
@@ 4.4 Ablation Study
4.5 Hyper-parameter Analysis
This section introduces the influence of the threshold parameter $\delta$ and the lambda parameter $\lambda$ on HiGNN. The $\delta$ modulates the similarity threshold of heterophilous information in constructing $A^{\prime}$ and the $\lambda$ balances the significance of the new channel within the HiGNN framework. As shown in Fig. 4a a smaller $\lambda$ leads to a decrease in model performance for both datasets, indicating the necessity of introducing heterophilous information. An overlarge $\lambda$ also leads to the degradation of the performance, especially for Chameleon. This implies the optimal $\lambda$ should be selected from $0.1$ to $1$. For the threshold $\delta$ as shown in Fig. 4b, the performance of HiGNN becomes better with the increase of $\delta$. This shows a larger $\delta$ could capture more efficient connections between semantically similar nodes, which correlates with the Theorem 0.A.1.
(a) Influence of Lambda $\lambda$
(b) Influence of Threshold $\delta$
@@ 4.5 Hyper-parameter Analysis
5 Related Work
5.1 Graph Neural Networks
Graph Neural Networks are effective for handling graph-structured data[24, 5]. These networks capture the dependencies between nodes through a message-passing mechanism. As a representative of GNNs, Graph Convolutional Networks (GCNs)[10] performs convolution on the spectral domain to aggregate information from neighbors, capturing the relationships between nodes in graphs. Graph Attention Networks (GATs)[17] specify different weights to different neighbors with attention mechanisms, enabling the model to focus on the most important information. For inductive representation learning, GraphSage[9] generates low-dimensional vectors by learning a function that aggregates neighbors. These GNNs operate effectively under the assumption of homophily, which posits that connected nodes tend to exhibit similarity. However, these methods experience significant performance degradation in heterophilous graphs, where connected nodes are more likely to have different labels.[25].
5.2 Graph Homophily
@@
6 Conclusion
This paper introduces the concept of heterophilous information as the distribution of node neighbors and proposes a practical solution to effectively leverage the rich semantics present in graphs. The new graph structure in our proposed HiGNN improves the connectivity between nodes with identical labels, thereby accommodating datasets with varying homophily, as shown by our theoretical and empirical analysis. Our experiments show the superiority of HiGNN in handling both homophilous and heterophilous graphs. It is worth noting that while HiGNN enhances the graph learning, it depends on the prior estimation of all node labels. This dependency not only hinges on the quality of label estimation but also increases computational complexity. In our future work, we aim to explore more efficient representations of heterophilous information that strike a balance between computational demands and performance.
Appendix 0.A Proof of Theorem 1
Theorem 0.A.1
The homophily degree $\hat{h}$ of $A^{\prime}$ is directly proportional to the threshold $\delta$, inversely proportional to the variance $\sigma^{2}$ of sampling noise, and directly proportional to the distance between the original homophily degree and reciprocal of c:
$Proof.$ We first make the following assumptions:
  • The number of nodes is balanced across each class.
Following the assumptions, we ensure the original homophily degree is $h$. Besides, we can get the heterophilous information of node $u$ as
$\mathcal{H}_{u}=[p_{1},p_{2},\cdots,p_{c}],where\ p_{i}=\begin{cases}h+\epsilon_{u,i}&\text{if }i=y_{u}\\ \frac{1-h}{c-1}+\epsilon_{u,i}&\text{if }i\neq y_{u}\end{cases}$ (11)
$\begin{split}\boldsymbol{S^{+}}&=\cos(\mathcal{H}_{u},\mathcal{H}_{v})\\ &=\frac{1}{\mathcal{H}_{norm}}\biggl{(}(h+\epsilon_{v,y_{v}})(h+\epsilon_{u,y_{u}})+\sum_{\begin{subarray}{c}i=1,\\ i\neq y_{u}\end{subarray}}^{c}(\frac{1-h}{c-1}+\epsilon_{u,i})(\frac{1-h}{c-1}+\epsilon_{v,i})\biggl{)}\\ \boldsymbol{S^{-}}&=\cos(\mathcal{H}_{u},\mathcal{H}_{v})\\ &=\frac{1}{\mathcal{H}_{norm}}\biggl{(}(h+\epsilon_{u,y_{u}})(\frac{1-h}{c-1}+\epsilon_{v,y_{v}})+(h+\epsilon_{v,y_{v}})(\frac{1-h}{c-1}+\epsilon_{u,y_{u}})\biggl{.}\\ \biggl{.}&+\sum_{\begin{subarray}{c}i=1,\\ i\neq y_{u},i\neq y_{v}\end{subarray}}^{c}(\frac{1-h}{c-1}+\epsilon_{u,i})(\frac{1-h}{c-1}+\epsilon_{v,i})\biggl{)}\end{split}$ (12)
[…]
where $\epsilon^{+}=h(\epsilon_{u,y_{u}}+\epsilon_{v,y_{v}})+\sum_{\begin{subarray}{c}i=1,i\neq y_{u}\end{subarray}}^{c}\frac{1-h}{c-1}(\epsilon_{u,i}+\epsilon_{v,i})\sim N(0,(h^{2}+\frac{(1-h)^{2}}{c-1})2\sigma^{2})$.
Then we can get the probability of homophilous edges as
Let $t_{+}=\frac{h^{2}+\frac{(1-h)^{2}}{c-1}-\delta\mathcal{H}_{norm}}{(h^{2}+\frac{(1-h)^{2}}{c-1})^{\frac{1}{2}}\sqrt{2}\sigma}$, we have $P(\boldsymbol{S^{+}}>\delta)=\Phi(t_{+})$.
Similarly, for $\boldsymbol{S^{-}}$, we follow the same process as $\boldsymbol{S^{+}}$ and we can get $P(\boldsymbol{S^{-}}>\delta)=\Phi(t_{-})$, where $t_{-}=\frac{2h\frac{1-h}{c-1}+(C-2)(\frac{1-h}{c-1})^{2}-\delta\mathcal{H}_{norm}}{(h^{2}+\frac{(1-h)^{2}}{c-1})^{\frac{1}{2}}\sqrt{2}\sigma}$.
From the partial derivative results, we can get Theorem 0.A.1 proved.

### v2

@@ IV-B Performance
Table III shows a comparison of HiGNN with other baselines. All methods are evaluated with node classification tasks, where both the mean accuracy and standard deviation are reported based on 10 public random splits. The average rank for each method is calculated across 9 datasets, with a lower rank indicating better performance. For each dataset, the best-performing methods are highlighted in bold and the runner-up methods are underlined. For methods that augment general GNN baselines with heterophilous information (denoted as +Hi), we halve the embedding space to control the total number of parameters and time complexity, ensuring a fair comparison. We use OOM to denote out-of-memory errors
@@
First, our proposed HiGNN generally outperforms all the other methods, achieving the lowest average rank, indicating the effectiveness of incorporating heterophilious information into graph learning. HiGNN enhances the performance on both the homophilous and heterophilous datasets, corroborating the explanation provided in Sect. III-B that the new matrix constructed by heterophilous information improves graph learning at different homophily levels. Besides, the effectiveness of heterophilous information is also shown in the improvement on the general GNNs baselines where the performance is improved on most of the datasets. Second, compared with general GNN baselines such as GCN and GAT, MLP performs poorly on homophilous graphs but still demonstrates strong performance on heterophilous graphs. This validates the assertion that heterophily impedes graph convolution in graph learning. Then for the methods that aggregate embeddings of k-hop neighbors during the message passing, Mixhop, H2GCN, GGCN, CPGCN, and GPR-GNN show their improvement with respect to GCN, but they still have a large gap with HiGNN and GloGNN that exploit global similar nodes. This indicates that for heterophilous graphs, focusing on k-hop neighbors is not as sufficient as constructing new graph structures that capture neighbors with similar semantics. Note that although GloGNN delivers the best results on 3 datasets, it requires the calibrated adjustment of many parameters444 Noted that some baseline results presented in Table3 maybe different from those in the original papers on some datasets. The performance of GloGNN[24] reported there is different as in their original paper because we use the preprocessed heterophilous datasets from [32]. In contrast, our method is more straightforward and flexible, requiring fewer parameter tunings.
TABLE III: Node classification performance of HiGNN
@@
IV-C Improvement in Homophily Degree
To validate that the new graph structure $A^{\prime}$ with heterophilous information in HiGNN improves homophily, we measure the homophily degree for both the original graph structure $A$ and the new graph structure $A^{\prime}$. Specifically, we compute homophily degrees across 9 datasets on 10 random splits and then report the mean and standard deviation. As depicted in Fig. 4, the newly constructed matrix improves the homophily degree on all datasets, particularly for heterophilous datasets. This suggests that heterophilous information is helpful in grouping semantically similar nodes. It is worth noting that the estimation of heterophilous information is based on the predicted labels instead of the true labels, demonstrating the estimation is error-tolerant as suggested in Sect. III-C.
@@ IV-C Improvement in Homophily Degree
IV-D Ablation Study
To demonstrate the performance improvement of HiGNN achieved through the incorporation of heterophilous information and the efficacy of the fusion strategy employed in HiGNN, we undertake an ablation study. As shown in Table IV, for HiGNN, we examine the influence of channels by removing the new channel (w/o $A^{\prime}$), removing the original channel (w/o $A$), or fusing two adjacency matrices $A$ and $A^{\prime}$ in one channel ($A$+$A^{\prime}$). The results show that the performance of HiGNN drops when either the new channel or the original channel is removed. This suggests that optimal performance is attained by considering different aspects of these two channels. Furthermore, the early fusion of two adjacency matrices ($A$+$A^{\prime}$) underperforms compared to the late fusion in HiGNN, indicating the late fusion strategy allows the model to fully capture the unique contributions of each channel as suggested in Sect. III-C.
TABLE IV: Ablation study on HiGNN
@@ IV-D Ablation Study
IV-E Hyper-parameter Analysis
This section introduces the influence of the threshold parameter $\delta$ and the lambda parameter $\lambda$ on HiGNN. The $\delta$ modulates the similarity threshold of heterophilous information in constructing $A^{\prime}$ and the $\lambda$ balances the significance of the new channel within the HiGNN framework. As shown in Fig. 5a a smaller $\lambda$ leads to a decrease in model performance for both datasets, indicating the necessity of introducing heterophilous information. An overlarge $\lambda$ also leads to the degradation of the performance, especially for Chameleon. This implies the optimal $\lambda$ should be selected from $0.1$ to $1$. For the threshold $\delta$ as shown in Fig. 5b, the performance of HiGNN becomes better with the increase of $\delta$. This shows a larger $\delta$ could capture more efficient connections between semantically similar nodes, which correlates with the Theorem III.1.
@@ IV-E Hyper-parameter Analysis
In addition, we report the hyper-parameters for HiGNN. As shown in Table V, HiGNN is not sensitive to most of the basic hyper-parameters including learning rate, dropout size, lambda, number of layer. This indicates stable performance of HiGNN in general GNN settings. Regarding the threshold $\delta$, the best results are achieved with high values, suggesting that a higher similarity of heterophilous information is always preferable. As for the lambda $\lambda$, the best results vary across different datasets, indicating that an optimal lambda $\lambda$ should be selected for each graph.
TABLE V: Hyper-parameters for HiGNN
Hyper-parameter Cora Citeseer Pubmed Chameleon Cornell Film Squirrel Texas Wisconsin
Learning Rate $10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-2}$
Dropout Rate 0.4 0.7 0.4 0.7 0.6 0.7 0.6 0.7 0.7
Embedding Size 128 256 128 256 128 64 256 128 256
Weight Decay $5\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$ $5\times 10^{-5}$
Lambda 1 1 $10^{-2}$ $10^{-4}$ $5\times 10^{-2}$ 1 $5\times 10^{-2}$ $10^{-2}$ 0.01
Layer Number 2 2 2 2 2 2 2 2 2
Threshold 0.8 0.9 1.0 0.8 1.0 0.8 0.8 0.9 1.0
V Related Work
V-A Graph Neural Networks
Graph Neural Networks are effective for handling graph-structured data[33, 34]. These networks capture the dependencies between nodes through a message-passing mechanism. As a representative of GNNs, Graph Convolutional Networks (GCNs)[3] perform convolution on the spectral domain to aggregate information from neighbors, capturing the relationships between nodes in graphs. Graph Attention Networks (GATs)[4] specify different weights to different neighbors with attention mechanisms, enabling the model to focus on the most important information. For inductive representation learning, GraphSage[5] generates low-dimensional vectors by learning a function that aggregates neighbors. These GNNs operate effectively under the assumption of homophily, which posits that connected nodes tend to exhibit similarity. However, these methods experience significant performance degradation in heterophilous graphs, where connected nodes are more likely to have different labels[7].
V-B Graph Homophily
@@
VI Conclusion
This paper introduces the concept of heterophilous information as the distribution of node neighbors and proposes a practical solution to effectively leverage the rich semantics present in graphs. In other words, the heterophily graph structure is regarded as semantic information rather than noise information. HiGNN is proposed to make good use of the semantics to improve the connectivity between nodes with identical labels, thereby accommodating datasets with varying homophily, as shown by our theoretical and empirical analysis. Our experiments show the superiority of HiGNN in handling both homophilous and heterophilous graphs. It is worth noting that while HiGNN enhances the graph learning, it depends on the prior estimation of all node labels. In our future work, we aim to explore more efficient representations of heterophilous information that strike a balance between computational demands and performance.